Lubricating oil worth 28 cents/quart is to be mixed with oil worth 33 cents/quart to make up 45 quarts of a mixture to sell at 30 cents/quart. What volume of each grade should be taken?
Found 2 solutions by josgarithmetic, TimothyLamb:
Answer by josgarithmetic(39617)

This kind of mixture question is very frequently asked, so this lesson is directed for this type:
http://www.algebra.com/my/mixture-price-two-part-both-parts-unknown.lesson?content_action=show_dev

Answer by TimothyLamb(4379)

x = quarts of 28 cent oil
y = quarts of 33 cent oil
---
x + y = 45
28x + 33y = 30*45
---
put the system of linear equations into standard form
---
x + y = 45
28x + 33y = 1350
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = quarts of 28 cent oil = 27
y = quarts of 33 cent oil = 18
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---